Partial bibliographic database for the ISO Legacy Conference

A. Coulais

April 30, 2002

References

LC_Beintema [1] D. A. Beintema and A. Salama, Beam profile analysis of SWS, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E40–+.

LC_Burgdorf [2] M. Burgdorf and A. Harwood, Radiation effects on LWS detectors and deglitching of LWS data, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E68–+.

LC_Caux [3] E. Caux, Transient effects correction for LWS detectors, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E47–+.

LC_Claret [4] A. Claret, Understanding and modelling glitch effects in ISOCAM, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E38–+.

LC_Coulais_GBT [5] A. Coulais and A. Abergel, The importance of ground based tests for space experiments using low background IR photodetectors, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E45–+.

LC_Coulais_Sources [6] A. Coulais and B. Fouks, Current status of modelization of point source transient response for LW-ISOCAM, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E73–+.

LC_Coulais_Models [7] A. Coulais and B. Fouks, Non–linear transient models for ir low-background detector, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E72–+.

LC_Coulais_PHT [8] A. Coulais, J. Sée, J. . Giovannelli, B. Stepnik, F. Balleux, G. Lagache, and A. Abergel, Transient correction for ISOPHOT C-100 detector, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E71–+.

LC_Burgo [9] C. del Burgo, P. Ábrahám, and U. Klaas, A new analysis of the dark current behaviour of the ISOPHOT detectors, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E74–+.

LC_Fouks [10] B. Fouks, Physical approach to the problems of space low-background detectors, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E43–+.

LC_Heras [11] A. M. Heras, E. Wieprecht, P. Nieminen, H. Feuchtgruber, F. Lahuis, K. Leech, R. Lorente, P. W. Morris, A. Salama, and B. Vandenbussche, Summary of the SWS detector radiation effects, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E37–+.

LC_Herbstmeier [12] U. Herbstmeier, P. Abrahám, P. Héraudeau, U. Klaas, R. J. Laureijs, T. Müller, and B. Schulz, Beam profiles of ISOPHOT detectors, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E41–+.

LC_Kaneda [13] H. Kaneda, M. Akazaki, Y. Okamura, T. Nakagawa, and I. Yamamura, Laboratory measurements of transient behavior of stressed ge:ga detectors, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E78–+.

LC_Kester [14] D. Kester, D. A. Beintema, and D. Lutz, SWS fringes and models, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E79–+.

LC_Kester_Fouks [15] D. Kester, B. Fouks, and F. Lahuis, SWS transient effects: Models and corrections, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E46–+.

LC_Lahuis [16] F. Lahuis, H. Feuchtgruber, H. Golstein, D. Kester, W. Luinge, and E. Wieprecht, Large SWS signal capture – from flux to signal –, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E82–+.

LC_Leeks [17] S. J. Leeks, B. M. Swinyard, T. L. Lim, and P. E. Clegg, The in-orbit performance of the LWS detectors, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E48–+.

LC_Lemke_Stray [18] D. Lemke, U. Klaas, and T. Kranz, Straylight in ISOPHOT ?, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E42–+.

LC_Lim_prelaunch [19] T. L. Lim, LWS pre-launch calibration, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E8–+.

LC_Lim_PHT [20] T. L. Lim, D. Lemke, and J. Wolf, The pre-flight calibration of the ISOPHOT instrument, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E9–+.

LC_Lim_Cal_LWS [21] T. L. Lim, B. M. Swinyard, M. Burgdorf, E. Caux, G. R. Davis, J. D. Ewart, A. M. di Giorgio, M. J. Griffin, C. Gry, A. S. Harwood, S. J. Leeks, S. Molinari, S. Pezzuto, S. D. Sidher, D. Texier, E. Tommasi, and N. R. Trams, LWS calibration strategy, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E3–+.

LC_Lloyd_franges [22] C. Lloyd, Fringing and spectrum fracturing in ISO LWS observations, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E40–+.

LC_Lloyd_beam [23] C. Lloyd, The ISO LWS beam profile, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E40–+.

LC_Nieminen [24] P. Nieminen, The in-orbit radiation environment and its effects on space-borne instrumentation, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E36–+.

LC_Rodighiero [25] G. Rodighiero and C. Lari, Total and ramp dependent internal linearity calibration of phot-c data, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E56–+.

LC_Salama [26] A. Salama, Spacecraft: Telescope and AOCS (pointing), in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E6–+.

LC_Shipman [27] R. F. Shipman, P. W. Morris, D. A. Beintema, D. Boxhoorn, H. Feuchtgruber, A. M. Heras, R. Huygen, D. Kester, F. Lahuis, K. Leech, R. Lorente, D. Lutz, P. Roelfsema, A. Salama, S. G. Schaeidt, E. Valentijn, B. Vandenbussche, and E. Wieprecht, SWS inflight calibration, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E18–+.

LC_Swinyard [28] B. Swinyard and T. Lim, Ground based calibration of the LWS, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001.

LC_Wilke [29] K. Wilke, U. Grözinger, U. Klaas, and D. Lemke, In-orbit curing procedures for ISOPHOT detectors, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E49–+.

LC_Yamamura [30] I. Yamamura and The ASTRO-F Team, Astro-f: Mission overview and the calibration strategy, in The calibration legacy of the ISO Mission’, proc. of a conference held Feb 5-9, 2001. Edited by L. Metcalfe and M. F. K. Kessler. To be published by ESA , Volume 481., 2001, pp. E61–+.